Masss of an object is measured from a faulty beam balance (mass of pans are not equal). It is 60 kg when the object is left pan and 70 Kg when the object is at right pan.Find the real mass of the object.

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

let masses of left pan right, pan and object be m,M and mo respectively

60+M=m+mo

70+m=M+mo

adding the above two equations, weget

130+m+M=m+M+2mo

therefore mass of object=65kg